OpenStudy (anonymous):

x^4 - 5x^2 -14 = 0 now substitute u = x^2 to solve for x

OpenStudy (nurali):

X^4 - 5X^2 -14= 0 Let U = X^2. This a "dummy variable" substitution. We get U^2 - 5U -14 = 0. We apply the quadratic to get U = (5 +- sqrt(25+56))/2 U = (5 +- sqrt(81))/2. We resubstitute X^2 for U and get X^2 = (5 +- sqrt(81))/2. X = +-sqrt((5 +-9)/2). X = +-sqrt((5 +9)/2) or X = +-sqrt((5 -9)/2) X = +-sqrt((14)/2) or X = +-sqrt((-4)/2)These are both real solutions.

OpenStudy (anonymous):

Nurali you are a great help thank you

OpenStudy (nurali):

Anytime.

OpenStudy (nurali):

Welcome to Openstudy

OpenStudy (anonymous):

Thank you

OpenStudy (nurali):

Anytime.